# Gibbs states and Brownian models for coexisting haze and cloud droplets

### v1

@@ RESULTS
In a developing cloud, turbulent updrafts and mixing with its neighboring air provoke supersaturation fluctuations that naturally affect droplet growth and, consequently, the collective droplet size spectra Grabowski and Abade (2017); Abade, Grabowski, and Pawlowska (2018). Stochastic models for the velocity fluctuations and the supersaturation fields have been considered in previous studies and have shown various degrees of relevance Sardina et al. (2015); Siewert, Bec, and Krstulovic (2017). For a cloud parcel subject to turbulent updrafts, the authors in Sardina et al. (2015) prove that activated droplet size variance increases, at short times, proportionally to the square-root of time, analogous to Einstein’s diffusion formula. Their analytical findings are contrasted with direct numerical simulations of turbulence obeying the incompressible Navier-Stokes equation. In spite of this diffusion result with diverging variance, in Siewert, Bec, and Krstulovic (2017), the authors find a statistical steady state for droplet size, with DSDs characterized by exponential tails; see their Eq. (4.3). Noteworthy is that to achieve convergence to a steady state, a substantial proportion of the droplets must reach evaporation, possibly if, on average, they experience negative supersaturations.
Supersaturation being the driving parameter, we assume that it evolves around a given value $\lambda$ with fluctuations encoded by a stochastic term possibly depending on the particle size. For the sake of generality, we aim at modeling the inhomogeneous effects of supersaturation fluctuations on droplets of different size. Indeed, this size-dependent fluctuations appear, for instance, when taking the adiabatic limit of fast supersaturation fluctuations, as demonstrated in (Siewert, Bec, and Krstulovic, 2017, Eq. (4.3)). Hence, incorporating different supersaturation sources and sinks alter the expression for the noise law affecting droplet growth. Such noise inhomogeneity is here motivated from the fact that small fluctuations in droplet radius incur greater fluctuations in supersaturation when the particles are bigger. Contrary to haze particles, which are at a stable thermodynamic equilibrium and variations in their radius provoke smaller alterations in the supersaturation budget.
@@
where $W_{t}$ is a Brownian motion (and $\dot{W}_{t}$ is Gaussian and white).
The function $V_{\lambda}$ is the potential function that collects the nonlinear, deterministic effects in the model. Thus, the solutions $X(t)$ to Eq. (8) can be regarded as Brownian particles embedded in the potential $V_{\lambda}$ defined in Eq. (8). The equilibria of Eq. (3) correspond exactly to the local critical points of the potential function—see Fig. 1c)—and their stability is determined by the local curvature of $V_{\lambda}$ at those points, see e.g. Berglund (2011); Pavliotis (2014). Note that Eq. (8) can always be recast into a form in which the stochastic disturbance term becomes $X$-independent, at the expense of changing the potential function via the Lamperti transformation (Pavliotis, 2014, §3.6). This trick is recalled in Methods as it is also part of our model’s analysis below.
For a family of monodisperse droplets in the same cloud volume, each one experiences different supersaturation fluctuations, albeit with the same mean and variance, as encapsulated by the term $\sigma(X)\dot{W}_{t}$ in Eq. (8). When each supersaturation’s stochastic realization is averaged over a large number of particles, the Fokker-Planck Equation (FPE) associated with Eq. (8) provides the DSD at a given time $t$; see e.g. Pavliotis (2014); Risken (1989). The stationary solution to the FPE is then given analytically and known as the Gibbs state (Pavliotis, 2014, Eq. (4.35)):
@@
where $Z_{\lambda}$ is a normalizing constant. This distribution is a generalization of what is found in (McGraw and Liu, 2006, Eq. (6)). Such equation yields a Weibull distribution for droplet radius $r$, without considering curvature and chemistry effects, i.e. Köhler theory, sink functions or state-dependent noise. One can, thus, derive the Weibull distribution $\rho_{W}$ for the radius $r$ by setting $f=g\equiv 0$ and $\sigma$ constant, where the resulting DSD is integrable only if there is mean subsaturation, $\lambda<0$:
@@
where $\gamma$ is a parameter proportional to $\lambda$, droplet number concentration and liquid water fraction; see (McGraw and Liu, 2006, Eq. (8)). In comparison, $\rho_{\lambda}$ in Eq. (9) can be integrable if $\lambda$ is positive, provided that the sink function $g$ has suitable decay properties detailed in Proposition 2 found in The confining potential in Methods. For instance considering sedimentation, or any size-independent removal mechanism, would provide an integrable DSD; see Krueger (2020) and (Chandrakar et al., 2020, Sections 2.1, 2.2 and 2.3), where the resulting DSDs are an explicit evaluation of $\rho_{\lambda}$ in Eq. (9).
For a broad class of $f$, $g$ and $\sigma$, the Gibbs state $\rho_{\lambda}$ exhibits multimodality intimately related to shape of the potential $V_{\lambda}$. More precisely, the Gibbs state’s modes, the local maxima of $\rho_{\lambda}$, are found by solving the following equation:
@@
See Noise-induced metastability in Methods for the derivation of this equation. This fundamental equation teaches us how the Gibbs state modes result from the interaction between the noise term and the potential function. The Gibbs state’s modes correspond to metastable states for which the random fluctuation term, $\sigma(X)\dot{W}_{t}$, in Eq. (8) acts as a source for the particles to experience transitions from a statistically typical droplet size to another one; see schematic of Fig. 3.
When the noise is state-independent, i.e. when $\sigma(X)=\sqrt{2\varepsilon}$, then $\sigma^{\prime}(X)=0$ and the metastable states are directly determined by the zeros of $-V_{\lambda}^{\prime}$. In this case, for Eq. (6), the Gibbs states are bimodal when $\lambda_{c}<\lambda<\lambda_{h}$, and unimodal when there is just one underlying stable equilibrium, i.e. for either $\lambda<\lambda_{c}$ or $\lambda>\lambda_{h}$; see Fig. 2. When $\lambda_{c}<\lambda<\lambda_{h}$, there is a non-vanishing probability of encountering haze or activated droplets; see Fig. 3. This property is consistent with realistic situations in which turbulent supersaturation fluctuations make haze susceptible of activating even if on average the Köhler’s threshold is not attained Prabhakaran et al. (2020).
Figure 3: Cartoon of haze-to-droplet stochastic transitions. Here $\lambda$ is kept fixed and the situation depicted corresponds to $\lambda$ in $(\lambda_{c},\lambda_{h})$ where coexist two stable states (haze particle and activated droplet); see Fig. 2.
Our modeling allows for other useful analytic insights. For instance, the expected residence times at the haze or activated state can be precisely determined. They are indeed known to relate to the potential barrier, $V_{\lambda}(X_{u})-V_{\lambda}(X_{h})$ (see Fig. 1c), according to the Kramers’ time formulas Kramers (1940); Berglund (2011). The mean activation time $\tau_{act}$ is then given by (Matkowsky and Schuss, 1981, Eq. (2.19))
@@
Therefore, the closer $\lambda$ is to the critical value $\lambda_{h}$ from below, the closer $X_{u}$ is to a collision with $X_{h}$ (see Fig. 2) and the shallower the potential barrier becomes, resulting into more common particle activation. Similarly, deactivation of activated droplets can take place due to fluctuations in supersaturation with a mean deactivation time obtained by (12) in which $X_{h}$ is replaced by $X_{c}$.
@@
Figure 4: Rationale of an hysteresis path. As $\lambda$ varies, the potential landscape changes and the stochastic disturbances can help trigger the critical transition as one approaches the tipping points, here the saddle-node bifurcation points.
@@
Figure 5: Panels a), b) and c): The thin gray line indicates the hysteresis loops for different rates of change of supersaturation $\dot{\lambda}$ as indicated in the title. The noise variance is constant for these three panels as indicated in the legend of Panel a). Panels d), e) and f): The thin gray lines indicate the hysteresis loops for different noise variances as indicated in the titles. The rate of change of supersaturation $\dot{\lambda}$ is constant for these three panels as indicated in the legend of Panel d). For all panels, the cyan line is the M-Köhler curve associated with Eq. (6) and $\beta(N)=3.6\times 10^{-2}\mathrm{s}^{-3/2}$.
@@
The hysteresis behavior pointed out at the end of Multistable Köhler curves is now analyzed when the supersaturation parameter $\lambda$ is allowed to drift through the critical values $\lambda_{c}$ and $\lambda_{h}$, slower compared to the stochastic fluctuations. Such relatively slower variations in $\lambda$ are associated with buoyancy-driven supersaturation changes. Supersaturation $\lambda$ increases as the buoyant parcel rises until the activation threshold $\lambda_{h}$ is attained and some of the contained haze is activated. When the parcel ceases to be buoyant or reaches a temperature inversion, supersaturation decreases to $\lambda_{c}$, below which the whole family of droplets regenerates into haze. During this buoyancy-cycle, the population of particles faces a continuous change in the potential $V_{\lambda}$, as illustrated in Fig. 4, for increasing and decreasing supersaturation levels. It is there shown the asymmetry of activation. Indeed, when a haze particle undergoes activation— starting at point A and following the red, wobbly line—, its deactivation will occur at much lower supersaturation levels— as seen when following the blue wobbly line that ends in point B—. Similar cycles have also been reported in the absence of aerosol curvature and chemistry effects, and uniquely due to noisy supersaturation fluctuations around a subsaturated mean; see (Siewert, Bec, and Krstulovic, 2017, Fig. 4).
@@
Within the present framework, it remains to investigate the effects of (i) different rates of changes in supersaturation and (ii) variations in the magnitude of the fast supersaturation fluctuations as encapsulated by the noise term. The hysteresis phenomenon under the presence of noise has been investigated in previous work Berglund and Gentz (2002), where it is found that a large (additive) noise variance $\sigma$, on average, reduces the area of hysteresis loops. In our context, this implies that droplet activation can occur at much lower supersaturation values compared to $\lambda_{h}$. To numerically examine these ideas, we solve the Eq. (8) associated with the sink term $g(X)=-\beta(N)X^{3/2}$ used in Eq. (6), for specific noise variances and mean supersaturation $\lambda$ ranging from $0.06\%$ to $0.11\%$ with different linear rate $\dot{\lambda}$.
@@
The numerical experiments regarding hysteresis are shown in Fig. 5, where Eq. (6) was integrated using an Euler-Maruyama scheme with a time step of $10^{-2}\mathrm{s}$. In Fig. 5a), b) and c), are shown hysteresis paths solving Eq. (6) for three rates of change $\dot{\lambda}$. As expected, when the rate of change in supersaturation is not slow enough, it can prevent droplet activation— see Fig. 5a)—, whereas sufficiently slow, adiabatic changes enforce the droplet size to follow the M-Köhler curve such as shown in Fig. 5c). As a haze particle experiences an increase in supersaturation— resulting from an adiabatic cooling of a cloud parcel— this will grow by condensation following the Köhler curve, almost regardless of how intense the cooling rate is, here indicated by $\dot{\lambda}$. This is an indication of the stability of the thermodynamic equilibrium of haze particles as opposed to cloud droplets, which are more sensitive the cooling rate $\dot{\lambda}$. Indeed, by comparing Figs. 5b) and c), we observe that only when the cooling rate is slowest do we observe a droplet growth attached to the Köhler curve.
Panels d), e) and f) of Figure 5 provide an illustration of results of Berglund and Gentz (2002) for Eq. (6). Here we show the effects of increasing turbulent noise intensity as a haze particle grows and activates in a slowly cooling air volume. In such case we are in the situation of Fig. 5c) where the particle’s size is dictated by the Köhler curve— modulo zero-mean fluctuations—. As noise variance is increased, the hysteresis loop is collapsed to the extent that droplet activation is expected as soon as $\lambda$ surpasses $\lambda_{c}$. Consequently, it is expected that slowly lifted humid parcels undergoing heaving turbulent updrafts— here encoded in the noise variance— can contain particles that activate at much lower supersaturations compared to that predicted by Köhler.
@@
Small, warm convective clouds are abundant in the atmosphere but hard to measure Koren et al. (2008). Such clouds are often characterized by weak updrafts and small supersaturation, allowing for a significant part of the aerosol to coexist as haze together with the activated part Altaratz et al. (2021); Hirsch et al. (2017). Cloud chambers are often used to study clouds in such thermodynamic conditions experimentally Prabhakaran et al. (2020); Chandrakar et al. (2016).
@@
In such work, the so-called Pi-chamber is seeded with monodisperse families of particles, and turbulent fluctuations are induced by changes in temperature, pressure, and aerosol injection rates. Disregarding transients, the system reaches a steady state in number concentration and size distributions are calculated spanning hazy and active particles. Depending on the supersaturation mean and variance, DSDs display different characteristics involving their multimodality and spread; see (Prabhakaran et al., 2020, Fig. 2). We show here that such experimental distributions are well approximated by the Gibbs states of our Brownian modelling framework once the sink and noise terms are appropriately designed. Our goal is to provide a qualitative and quantitative correspondence between the present theory and the experiments from Prabhakaran et al. (2020).
We first address the sink term design. Note that for a monodisperse family of particles embedded in a cloud parcel, the supersaturation budget must be included in the collective droplet growth Squires (1956); Korolev and Mazin (2003). Because the Pi-chamber is closed and in steady-state, the sink term affecting supersaturation is determined by water vapor condensation onto the particles. The mean sources of supersaturation are taken to be zero, apart from fluctuations encoded in $\sigma(X)$; see Eq. (8). Consequently, and to a first approximation, the initially prescribed supersaturation $\lambda$ is consumed proportionally to the radius of the droplets, so that it evolves according to:
$\lambda(t)\approx\lambda-\beta X(t)^{1/2}+\sigma(X)\dot{W}_{t},$ (13)
where the absorption parameter $\beta$ depends on the number of active droplets, and the effective diffusivity of water vapour in air Khain and Pinsky (2018). By substituting $\lambda$ in the (standard) Köhler equation (1), the new term $-\beta X(t)^{1/2}$ plays the role of a sink term in Eq. (3), so that $g(X)=-\beta X^{1/2}$.
Regarding the term $\sigma(X)\dot{W}_{t}$, the inhomogeneous effect of fluctuations in haze and cloud droplets becomes apparent by qualitatively examining the DSDs obtained in the Pi-chamber; see (Prabhakaran et al., 2020, Fig. 2). To account for such dependence on particle size, we propose the following stochastic parameterization:
@@
$\psi(X)=\tanh(2\kappa D(X-X^{\ast})).$
Here, the hyperbolic tangent is employed as a step function that allows for interpolating smoothly between the noise strengths that affect haze and activated particles, $\sigma_{1}$ and $\sigma_{2}$ respectively. Thus, we require that $\sigma_{2}>\sigma_{1}>0$ for the reason stated above. The value of the slope is chosen to be sufficiently large— i.e., $102D\kappa=$ 10— so that the hyperbolic tangent approximates a step function, while still exhibiting the attributes of a smooth function, more amenable for analysis and numerical simulations. The “ignition” parameter $X^{\ast}$ is chosen to be equal to $6.2\cdot 10^{-3}\mathrm{s}$ or, in diameter, $d^{\ast}=1.41\mu\mathrm{m}$, which yields a correct location of the haze modes, and it is in the order of the critical Köhler diameter which is $d_{K}=1.7\mu\mathrm{m}$. Such critical value is derived from Eq. (1), as explained in Köhler theory, for the curvature and solubility coefficients $A=1.4\cdot 10^{-3}\mu\mathrm{m}$ and $B=3.5\cdot 10^{-4}\mu\mathrm{m}^{3}$; see Prabhakaran et al. (2020).
@@
Figure 6: Panels a), b) and c): Köhler and M-Köhler curves (dashed and solid curves, respectively) for each supersaturation value considered in PI-Chamber Empirical Distribution vs Gibbs States (Cases I, II, and III). The parameters of the Köhler curve are $A=1.4\cdot 10^{-3}\mathrm{\mu m}$ and $B=3.5\cdot 10^{-4}\mathrm{\mu m^{3}}$ and the other parameters of model Eq. (15) are given in Table 2. The black dot denotes the location of the metastable state induced by the state-dependent noise; see Noise-induced metastability in Methods. Panels d), e) and f): Experimental DSDs from Prabhakaran et al. (2020) vs Gibbs states from Eq. (9). Both the data histograms and Gibbs states are normalized according to their integral. The insets show the transformed potentials (due to Lamperti transformation) explaining the shapes of the Gibbs states; see The Lamperti transformation in Methods.
We now confront the ability of our Brownian model (15) to approximate experimental DSDs by the Gibbs states formula (9). Three experimental setups are considered in that respect, following those of the Pi-chamber experiments from Prabhakaran et al. (2020) corresponding to different supersaturation mean and fluctuation properties reported in such paper. These are referred to as according to Prabhakaran et al. (2020): a) mean-dominated, b) fluctuation influenced and c) fluctuation dominated acivation.

### v2

@@ Results
In a developing cloud, turbulent updrafts and mixing with its neighboring air provoke supersaturation fluctuations that naturally affect droplet growth and, consequently, the collective droplet size spectra. In that respect, we refer to [DSS+12, SS17] for observational studies in stratocumulus clouds, and to [PS09, KRZC97] regarding the influence of temperature and pressure fluctuations in supersaturation distribution. Stochastic models for the velocity fluctuations and the supersaturation fields have been considered in previous studies and have shown various degrees of relevance [SPBC15, SBK17]. For a cloud parcel subject to turbulent updrafts, the authors in [SPBC15] prove that activated droplet size variance increases, at short times, proportionally to the square-root of time, analogous to Einstein’s diffusion formula. Their analytical findings are contrasted with direct numerical simulations of turbulence obeying the incompressible Navier-Stokes equation. In spite of this diffusion result with diverging variance, in [SBK17], the authors find a statistical steady state for droplet size, with DSDs characterized by exponential tails; see their Eq. (4.3). Noteworthy is that to achieve convergence to a steady state, a substantial proportion of the droplets must reach evaporation, possibly if, on average, they experience negative supersaturations.
Supersaturation being the driving parameter, we assume that it evolves around a given value $\lambda$ with fluctuations encoded by a stochastic term possibly depending on the particle size. The rationale behind this modeling assumption is that larger droplets are expected to have a greater impact on the supersaturation budget than the smaller ones, in a limited water vapor field. This is what we here referred to as inhomogeneous fluctuations, the inhomogeneous character being dependent on the droplets’ size. The type of inhomogeneity referred to here should not be confused thus with that caused by inhomogeneous mixing in clouds, occurring under large entrainment events [KKPF16].
@@
where $W_{t}$ is a Brownian motion (and $\dot{W}_{t}$ is Gaussian and white). An example of the $\sigma(X)$ function is given in Section "Pi-Chamber empirical distributions vs Gibbs states" below.
The function $V_{\lambda}$ is the potential function that collects the nonlinear, deterministic effects in the model. Thus, the solutions $X(t)$ to Eq. (8) can be regarded as Brownian particles embedded in the potential $V_{\lambda}$ defined in Eq. (8). The equilibria of the deterministic counterpart of Eq. (8) (i.e. Eq. (3)) correspond exactly to the local critical points of the potential function—local minima in Fig. 1c)—and their stability is determined by the local curvature of $V_{\lambda}$ at those points, see e.g. [Ber11, Pav14]. Note that Eq. (8) can always be recast into a form in which the stochastic disturbance term becomes $X$-independent, at the expense of changing the potential function via the Lamperti transformation; see Chapter 3.6 in [Pav14]. This trick is recalled in Section "Lamperti transformation" in Material and Methods as it is also part of our model’s analysis below.
For a family of monodisperse droplets in the same cloud volume, each one experiences different supersaturation fluctuations, albeit with the same mean and variance, as encapsulated by the term $\sigma(X)\dot{W}_{t}$ in Eq. (8). When each supersaturation’s stochastic realization is averaged over a large number of particles, the Fokker-Planck Equation (FPE) associated with Eq. (8) provides the DSD at a given time $t$; see e.g. [Pav14, Ris89]. The stationary solution to the FPE is then given analytically and known as the Gibbs state (see [JKO98] and Eq. (4.35) in [Pav14]):
@@
where $Z_{\lambda}$ is a normalizing constant. This distribution is a generalization of Weibull distribution found in Eq. (6) in [ML06] to explain diffusive broadening of the DSDs in turbulent clouds. Within our framework, it corresponds to Gibbs states without sink function and state-dependency in the noise, i.e. density distribution $\rho_{W}$ (for the radius $r$) obtained by setting $f=g\equiv 0$ and $\sigma$ constant. In this particular case, the resulting DSD is integrable only if there is mean subsaturation, $\lambda<0$ and is given by:
@@
where $\gamma$ is a parameter proportional to $\lambda$, droplet number concentration and liquid water fraction; see Eq. (8) in [ML06]. In comparison, $\rho_{\lambda}$ in Eq. (9) can be integrable if $\lambda$ is positive and subject to size-dependent fluctuations, provided that the sink function $g$ has suitable decay properties such as detailed in Proposition 2; see Section "The confining potential" in Material and Methods. We refer to [CSY+20] for other theoretical approaches to infer integrable DSDs for fluctuating supersaturation with size-dependent removal terms.
Our framework, for a broad class of functions $f$, $g$ and $\sigma$, allows for Gibbs state $\rho_{\lambda}$ that exhibits multimodality intimately related to shape of the potential $V_{\lambda}$. More precisely, the Gibbs state’s modes, the local maxima of $\rho_{\lambda}$, are found by solving the following equation:
@@
See Section "Noise-induced metastability" in Material and Methods for the derivation of this equation. This fundamental equation teaches us how the Gibbs state modes result from the interaction between the noise term and the potential function. The Gibbs state’s modes correspond to metastable states for which the random fluctuation term, $\sigma(X)\dot{W}_{t}$, in Eq. (8) acts as a source for the particles to experience transitions from a statistically typical droplet size to another one.
When the noise is state-independent, i.e. when $\sigma(X)=\sqrt{2\varepsilon}$, for some positive constant $\varepsilon>0$ (controlling the noise variance), then $\sigma^{\prime}(X)=0$ and the metastable states are directly determined by the zeros of $-V_{\lambda}^{\prime}$. In this case, for Eq. (6), the Gibbs states are bimodal when $\lambda_{c}<\lambda<\lambda_{h}$, and unimodal when there is just one underlying stable equilibrium, i.e. for either $\lambda<\lambda_{c}$ or $\lambda>\lambda_{h}$, as shown in Fig. 2, whereby the solid points denote the presence of equilibria, ranging from a single haze equilibrium to an activated one.
When $\lambda_{c}<\lambda<\lambda_{h}$, there is a notable chance of encountering haze or activated droplets. As depicted in Fig. 3, the red curve represents a potential energy landscape with two local minima similar to that shown in Fig. 1c). Stochastic fluctuations can then overcome the potential barrier between these two minima, leading to activation even when average supersaturation falls below the Köhler threshold. This phenomenon aligns with real-world scenarios where turbulent supersaturation variations can induce haze activation [PSK+20].
Figure 3. Cartoon of haze-to-droplet stochastic transitions. Here $\lambda$ is kept fixed and the situation depicted corresponds to $\lambda$ in $(\lambda_{c},\lambda_{h})$ where two stable states coexist (haze particle and activated droplet); see Fig. 2.
Our modeling framework allows for other useful analytic insights. For instance, the expected residence times at the haze or activated state can be precisely determined. They are indeed known to relate to the potential barrier, $V_{\lambda}(X_{u})-V_{\lambda}(X_{h})$ (see Fig. 1c), according to the Kramers’ time formulas [Kra40, Ber11]. The mean activation time $\tau_{act}$ is then given by Eq. (2.19) in [MS81]
@@
Therefore, the closer $\lambda$ is to the critical value $\lambda_{h}$ from below, the closer $X_{u}$ is to $X_{h}$ (see Fig. 2) and the shallower the potential barrier becomes, resulting into more common particle activation. Similarly, deactivation of activated droplets can take place due to fluctuations in supersaturation with a mean deactivation time obtained by (12) in which $X_{h}$ is replaced by $X_{c}$.
@@
Figure 4. Rationale of an hysteresis path. As $\lambda$ varies, the potential landscape changes and the stochastic disturbances can help trigger the critical transition as one approaches the tipping points, here the saddle-node bifurcation points.
@@
Figure 5. A landscape of hysteresis loops. Panels a), b) and c): The thin gray line indicates the hysteresis loops for different rates of change of supersaturation $\dot{\lambda}$ as indicated in the title. The noise variance is constant for these three panels as indicated in the legend of Panel a). Panels d), e) and f): The thin gray lines indicate the hysteresis loops for different noise variances as indicated in the titles. The rate of change of supersaturation $\dot{\lambda}$ is constant for these three panels as indicated in the legend of Panel d). For all panels, the cyan line is the M-Köhler curve associated with Eq. (6) where $\beta(N)=3.6\times 10^{-2}\mathrm{s}^{-3/2}$.
@@
The hysteresis behavior pointed out at the end of Section "Multistable Köhler curves" is now analyzed when the supersaturation parameter $\lambda$ is allowed to drift through the critical values $\lambda_{c}$ and $\lambda_{h}$, slower compared to the stochastic fluctuations. Such relatively slower variations in $\lambda$ are associated with buoyancy-driven supersaturation changes. Supersaturation $\lambda$ increases as the buoyant parcel rises until the activation threshold $\lambda_{h}$ is attained and some of the contained haze is activated. When the parcel ceases to be buoyant or reaches a temperature inversion, supersaturation decreases to $\lambda_{c}$, below which the whole family of droplets regenerates into haze.
During this buoyancy-cycle, the population of particles faces a continuous change in the potential $V_{\lambda}$, as illustrated in Fig. 4, for increasing and decreasing supersaturation levels. It is there shown the asymmetry of activation. Indeed, when a haze particle undergoes activation— starting at point A and following the red, wobbly line—, its deactivation will occur at much lower supersaturation levels— as seen when following the blue wobbly line that ends in point B—. Similar cycles have also been reported in the absence of aerosol curvature and chemistry effects, and uniquely due to noisy supersaturation fluctuations around a subsaturated mean; see e.g. Fig. 4 in [SBK17].
@@
Within the present framework, it remains to investigate the effects of (i) different rates of changes in supersaturation and (ii) variations in the magnitude of the fast supersaturation fluctuations as encapsulated by the noise term. The hysteresis phenomenon under the presence of noise has been investigated in previous work [BG02], where it is found that a large (additive) noise variance $\sigma$, on average, reduces the area of hysteresis loops. In our context, this implies that droplet activation can occur at much lower supersaturation values compared to $\lambda_{h}$. To numerically examine these ideas, we solve Eq. (8) associated with the sink term $g(X)=-\beta(N)X^{3/2}$ used in Eq. (6), for specific noise variances and mean supersaturation $\lambda$ ranging from $0.06\%$ to $0.11\%$ with different rate of change $\dot{\lambda}$. We refer to Section "Hysteresis path algorithm" in Material and Methods for numerical details.
Our model supports different hysteresis cycles as shown in Figure 5. The nature of these hysteresis loops is controlled by the rate of change of supersaturation $\dot{\lambda}$ and the intensity of the turbulent effects. Figure 5a) to Figure 5c) correspond to an increasingly faster response of droplet growth to supersaturation changes. Figure 5d) to Figure 5f) correspond to an increasing susceptibility to activation as the noise variance is increased, i.e. the turbulent effects are more pronounced while the rate of change $\dot{\lambda}$ is kept fixed.
@@
As expected, when the rate of change in supersaturation is not slow enough, it can prevent droplet activation— see Fig. 5a)—, whereas sufficiently slow, adiabatic changes enforce the droplet size to follow the M-Köhler curve such as shown in Fig. 5c). As a haze particle experiences an increase in supersaturation— resulting from an adiabatic cooling of a cloud parcel— this will grow by condensation following the Köhler curve, almost regardless of how intense the cooling rate is, here indicated by $\dot{\lambda}$. This is an indication of the stability of the thermodynamic equilibrium of haze particles as opposed to cloud droplets, which are more sensitive the cooling rate $\dot{\lambda}$. Indeed, by comparing Figs. 5b) and c), we observe that only when the cooling rate is slowest do we observe a droplet growth attached to the Köhler curve.
Panels d), e) and f) of Figure 5 provide for Eq. (6), illustrations of hysteresis cycles that are consistent with those analyzed for large-noise regimes in theoretical studies [BG02]. Compared to Panels a) and b) of Figure 5, the rate of change of supersaturation $\dot{\lambda}$ is smaller, resulting into sharper transitions to activation similar than those shown in Figure 5c). This combined with a larger turbulent noise intensity than in the cases of Figure 5a)-b)-c), may provoke transitions to occur before reaching the haze-to-droplet saddle-node bifurcation, i.e. particles that activate at much lower supersaturations compared to that predicted by Köhler. Physically it corresponds to haze particles that grow in a slowly cooling air volume embedded in a highly turbulent environment undergoing heaving turbulent updrafts. In some dramatic limit, the hysteresis loop may collapse such as shown in Figure 5f). This extreme situation is to put in contrast with the other extreme situation shown in Figure 5a) where activation is "unrealized" due to faster rate of change of supersaturation and weakly turbulent updraft.
@@
Small, warm convective clouds are abundant in the atmosphere but hard to measure [KOF+08]. Such clouds are often characterized by weak updrafts and small supersaturation, allowing for a substantial part of the aerosol to coexist as haze together with the activated part [AKA+21, HKA+17]. Cloud chambers are often used to study clouds in such thermodynamic conditions experimentally [PSK+20, CCC+16].
@@
In such work, the so-called Pi-chamber is seeded with monodisperse families of particles, and turbulent fluctuations are induced by changes in temperature, pressure, and aerosol injection rates. Disregarding transients, the system reaches a steady state in number concentration and DSDs are calculated spanning haze and active particles. Depending on the supersaturation mean and variance, DSDs display different characteristics involving their multimodality and spread; see Fig. 2 in [PSK+20]. We show here that such experimental distributions are well approximated by the Gibbs states of our Brownian modelling framework once the sink and noise terms are appropriately designed. Our goal is to provide a qualitative and quantitative correspondence between the present theory and the experiments from [PSK+20].
We first address the sink term design. Note that for a monodisperse family of particles embedded in a cloud parcel, the supersaturation budget must be included in the collective droplet growth [Squ56, KM03]. Because the Pi-chamber is in steady-state, the sink term affecting supersaturation is determined by water vapor condensation onto the particles, which boils down to assuming that supersaturation is consumed proportionally to the radius of the droplets. Thus a sink term of type $g(X)=-\beta X^{1/2}$ is chosen in Eq. (8), where the absorption parameter $\beta$ depends on the number of active droplets, and the effective diffusivity of water vapor in air [KP18]. It must be noted, however, that although in steady-state, the Pi-chamber is not a closed system. There is an intrinsic decorrelation timescale of supersaturation; see [CCC+16]. This timescale is usually determined by the coefficients in front of the deterministic terms of a (linear) stochastic Langevin equation governing the supersaturation evolution; see e.g. Eq. (2) in [CCC+16].
In our formalism, while not solving an explicit governing equation for supersaturation we can still determine a decorrelation timescale for supersaturation albeit in an indirect fashion. One first solves Eq. (8), records $X(t)$ and then form supersation time series according to:
$\lambda(t)=\lambda-\beta X(t)^{1/2}+\sigma(X(t))\dot{W}_{t}.$ (13)
Now the expectation of the stochastic term $\sigma(X)\dot{W}_{t}$ is zero due to the martingale property of Brownian motion and the independence of $\sigma(X(t))$ and $\dot{W}_{t}$. This is actually a fundamental property of stochastic differential equations driven by Brownian motion; see Section 4.2.6 in [Gar09]. Therefore the expected value of supersaturation has a decorrelation time given by that of $\lambda-\beta\mathbb{E}X(t)^{1/2}$, that is a function of the decorrelation time of the droplet radius. Thus the droplet size effects on supersaturation fluctuations encoded by the stochastic term $\sigma(X)\dot{W}_{t}$ affect the decorrelation time of the expected value of supersaturation in an indirect way, after solving Eq. (8) to get access to $X(t)$.
In the case of the Pi-chamber datasets analyzed below, we model the size-dependent effects on supersaturation fluctuations via the following functional:
@@
$\psi(X)=\tanh(2\kappa D(X-X^{\ast})),$
where we require that $\sigma_{2}>\sigma_{1}>0$ so that larger droplets have a bigger impact on supersaturation fluctuations than the smaller ones. The value of the slope is chosen to be sufficiently large — i.e., $102\kappa D=$ 10 — so that the hyperbolic tangent function, $\psi$, approximates a step function, while still exhibiting the attributes of a smooth function, more amenable for analysis and numerical simulations. The “ignition” parameter $X^{\ast}$ is chosen to be equal to $6.2\cdot 10^{-3}\mathrm{s}$ or, in diameter, $d^{\ast}=1.41\mu\mathrm{m}$, which yields a correct location of the haze mode, and it is in the order of the critical Köhler diameter which is $d_{K}=1.7\mu\mathrm{m}$. Such a critical value is derived from Eq. (1), as explained in Köhler theory, for the curvature and solubility coefficients $A=1.4\cdot 10^{-3}\mu\mathrm{m}$ and $B=3.5\cdot 10^{-4}\mu\mathrm{m}^{3}$; see [PSK+20].
@@
Figure 6. Pi-chamber experimental distributions vs Gibbs states, and underlying M-Köhler curves. Panels a), b) and c): Köhler and M-Köhler curves (dashed and solid curves, respectively) for each supersaturation value considered in Section "Pi-Chamber Empirical Distribution vs Gibbs States" (Cases I, II, and III). The parameters of the Köhler curve are $A=1.4\cdot 10^{-3}\mathrm{\mu m}$ and $B=3.5\cdot 10^{-4}\mathrm{\mu m^{3}}$ and the other parameters of model Eq. (15) are given in Table 2. The black dot denotes the location of the metastable state induced by the state-dependent noise; see Section "Noise-induced metastability" in Material and Methods. Panels d), e) and f): Experimental DSDs from [PSK+20] vs Gibbs states from Eq. (9). Both the data histograms and Gibbs states are normalized according to their integral. The insets show the transformed potentials (due to Lamperti transformation) explaining the shapes of the Gibbs states; see Section "The Lamperti transformation" in Material and Methods.
We now confront the ability of our Brownian model (15) to approximate experimental DSDs by the Gibbs states formula (9). Three experimental setups are considered in that respect, following those of the Pi-chamber experiments from [PSK+20] corresponding to different supersaturation mean and fluctuation properties reported in such paper. These are referred to as according to [PSK+20]: a) mean-dominated, b) fluctuation influenced and c) fluctuation dominated acivation.